### v1

# Scalable registration of single quantum emitters within solid immersion lenses through femtosecond laser writing

Abstract
Optically active silicon-vacancy (VSi) centers in silicon carbide (SiC) serve as qubits, interfacing spins via photons. This capability allows the encoding of photonic information within the spin state and facilitates on-demand readout, promising applications such as quantum memories. However, electron irradiation, a common technique for creating defects in SiC, lacks spatial selectivity, limiting scalability. We employed femtosecond laser writing within photonic structures to generate single (VSi) centers, registering them to photonic structures and enhancing optical collection efficiency by a factor of 4.5. Characterization of 28 laser-written defects centers in solid immersion lenses (SILs) showed distributions relative to the photonic structure’s center of 260 nm in the x-direction and 60 nm in the y-direction, with standard deviations of () nm and () nm, respectively. This method is scalable for developing integrated quantum devices using spin-photon interfaces in SiC.
I Introduction
Color centers in the solid state, such as optically-active point defects and impurities, are among the most prominent systems for quantum technology [1]. Spin-photon interfaces associated to color centers in diamond [2, 3, 4], silicon carbide (SiC) [5, 6, 7, 8, 9] and silicon (Si) [10, 11], along with single rare-earth ion dopants in crystals [12], have been used in pioneering demonstrations of long-distance quantum networks. The nitrogen-vacancy center in diamond, as well as other defects in SiC [13, 14, 15, 16] and hexagonal boron nitride [17, 18, 19, 20], are used in a variety of quantum sensing applications, ranging from probing condensed matter problems and fundamental physics, to healthcare [21].
Enhancing photon collection efficiency through tailored photonic structures is critical for optically-read-out spin qubits, as it directly impacts spin-photon interface efficiency in quantum networks and sensitivity in quantum sensing [22, 23]. This has been achieved for example with simple structures that minimize the effect of total internal reflection inside a high-index material, such as solid immersion lenses (SILs) [24, 25, 26, 27], or light guiding structures such as waveguides [28, 29, 30, 31]. The alternative is to employ a microcavity, such as a nanopillar [32, 33, 34, 35, 36] or photonic crystal [37, 38], to enhance light-matter interaction, maximize the fraction of emission into the coherent zero-phonon line, and increase optical extraction efficiency [39, 40].
A significant challenge in the scalable integration of quantum emitters into photonic structures is their registration into the area within the structure that provides the most substantial optical enhancement. Depending on the specific structure, this region can be on the order of about in SILs, down to a few tens of nanometers for resonant structures with small mode volumes, such as photonic crystal cavities. Two main approaches have registered single quantum emitters to photonic structures. The first consists of mapping the position of pre-existing emitters against an array of markers and then fabricating the structure concerning the same marker array [41, 42, 43]. This procedure provides accuracy down to a few nanometers [41] and enables pre-selecting emitters with optimal properties for the specific application, but it is extremely time-consuming and not easily scalable into large arrays.
A second possibility is to implant the ion species required to create the emitter into existing photonic structures, for example, through the use of focused ion beams [44, 45, 46, 47]. This technique is very effective, but can only be used to create shallow emitters (depths of m depth), as the lateral accuracy of implanted ion placement is similar in scale to the implantation depth due to increased straggle within the material. Furthermore, the higher ion energies required for deeper implantation typically induce more considerable lattice damage, which may degrade the quantum emitter properties [48, 49, 50].
Recently, the generation of quantum emitters by laser-writing has received increasing attention. This technique exploits highly energetic carriers, created either by below-bandgap illumination with a high-power femtosecond laser [51, 52, 53, 54] or above-bandgap illumination with lower laser power [55], to initiate an avalanche ionization process. Above-bandgap (UV) illumination has been used to create quantum emitters in SiC nanophotonic structures [55], and femtosecond NIR laser pulses have been utilized to generate ensembles of quantum emitters in diamond nano-cavities [45].
Here, we create single quantum emitters registered to SILs by direct below-bandgap femtosecond laser writing. Femtosecond laser writing enables us to create intrinsic point defects based on vacancies directly in the focal region of the SILs. Therefore, this method does not require any alignment to markers and can be fast and compatible with wafer-scale processing. The photonic structure itself can enhance the laser-writing field, potentially removing the need for any registration effort. By using the focusing effects of the lens, weaker laser pulse energies can be used to generate emitters. In contrast to ion implantation and UV illumination, direct below-bandgap femtosecond laser-writing enables the creation of single quantum emitters deeper into micron-scale structures, which are less sensitive to surface noise and typically feature better quantum coherence properties.
We focus on quantum emitters in SiC, a material that uniquely combines spin-photon interfaces [5, 6, 8] with long spin coherence times [56, 7] with the possibility of integrating photonic [57, 58, 59, 28] and microelectronic functionalities [60, 61]. Our approach readily adapts to alternative materials, such as diamond [24, 42, 62].
II Experiment details
The experiment was conducted using commercial 4H-SiC material (Xiamen PowerWay©), with substrate and epilayer thickness and respectively, and residual n-doping level of . The material was diced into chips, and arrays of hemispherical SILs with nominal radius were fabricated using the grayscale hard-mask lithography process set out in [26]. It is worth noting that this method provides a key combination of scalability to large arrays with minimal time and cost increase and control over the shape of high-aspect-ratio microstructures, as opposed to other techniques which are either slow and expensive (e.g. focused ion beam milling [63] and two-photon polymerization [64, 65]) or provide limited shape control and aspect ratio (e.g. resist reflow lithography [66, 67, 68, 69], self-assembled microspheres [70], and microscale screen printing [71]).
The ability to create large arrays of shape-controlled lenses enables systematic parameter characterization when laser writing into SILs, an essential step due to the beam-focusing effects of the lens. This effect means that the parameters from writing through a planar interface cannot be straightforwardly applied within SILs, as described in the following section.
II.1 Laser writing procedure
The laser writing process for generating quantum defects in the center of prefabricated SILs in 4H-SiC is illustrated in Fig. 1 a. The laser writing was performed using a regeneratively amplified Ti:Sapphire laser source, delivering 250 fs pulses at a wavelength of 790 nm and a repetition rate of 1 kHz (shown in Supplementary Figure S2). The laser was focused through a high-NA oil immersion objective lens (Olympus 60, 1.4 NA).
A single femtosecond laser pulse with adjustable pulse energy was directed to selected focal points within the 4H-SiC sample. The degree of induced lattice damage can be controlled by varying the pulse energy, both in the case of writing in the planar region and within SILs. For planar interfaces, the focal depth was set to 5 m, matching the radius of the hemispherical monolithic SILs (depicted by the red spots in Fig. 1 a). Focusing into high-index materials like 4H-SiC induces strong spherical aberrations, which were corrected using a spatial light modulator (SLM).
During fabrication, the laser was focused through the SIL to its central axis at a depth corresponding to the sample’s planar surface (represented by the blue spots in Fig. 1 a). This adjustment aimed to enable precise defect placement with minimal lattice perturbation and ensure optimal axial alignment between the SIL focus and laser induced defects.
The laser writing process was initially characterized by photoluminescence (PL) under 532 nm CW optical excitation (1 mW) in a home-built confocal microscope. In both SIL and planar configurations, the intensity of the PL decreases with a lower pulse energy, as shown in Fig. 1 b. For each pulse energy level, data were collected from five spatially separated sites (in the case of SILs, 5 different neighboring SILs) fabricated in a single column. All data points represent the arithmetic mean of intensities measured across these five equivalent fabrication points. Saturating behavior is observed in both cases, where the PL intensities detected from defect centers stabilize with respect to the writing power in the regime of high pulse energy.
Notably, significantly lower pulse energies are required for laser writing through the SIL interface compared to the planar interface. This is evidenced by the plotted curve shown in Fig. 1 b shifting to a lower pulse energy regime for the SIL interface (1.4 to 4.0 nJ) relative to the planar interface (4.0 to 20 nJ), highlighting the additional focusing effect provided by the SIL. Furthermore, the saturation count rate for the SIL interface reaches 600 kcps, compared to 300 kcps for the planar interface. The upward shift in the saturation curve indicates the enhanced light extraction efficiency achieved through the SIL interface.
II.2 Laser writing-induced defect generation mechanisms
Defect generation by ultrafast laser pulses in transparent materials is typically modeled by three mechanisms: multiphoton ionization (MPI), tunneling ionization (Zener breakdown), and avalanche ionization [72]. We take the model for generating the GR1 defect (neutral carbon vacancy) in diamond [73] as our starting point. The 250 fs laser pulses used here are shorter than the characteristic (nanosecond) timescales of thermal diffusion [74], allowing thermal effects to be neglected. The transition between MPI and tunneling ionization is described by the Keldysh parameter [72]. MPI dominates over tunneling ionization when
where kg is the effective mass of lattice electrons, and eV are the refractive index and bandgap of 4H-SiC and is the frequency of the photons. Substituting these values, the intensity threshold is calculated as . The pulse energy is then determined as , where the beam waist is nm (for the planar interface) and pulse duration is fs. Substituting these parameters for planar interface. For the SIL interface, the effective NA increases to around 2.6, which reduces the beam waist to nm, and the pulse energy is then calculated to be . The curvature of the SIL increased the effective NA , tightening the beam waist and intensifying the field strength in the focal volume, which ultimately lowers the pulse energy threshold.
For very high pulse energies, MPI does not dominate and additional energy may not create new defects due to saturation of the available defect sites. In contrast to the model in [73] for the GR1 center in diamond, we consider a saturation model alongside the power law:
where is the amplitude coefficient, is the power law exponent, and is the saturation parameter. Experimental results for the SIL interface yield and , while for the planar interface, and . The enhanced non-linearity in SIL compared to the planar interface may arise from a stronger Zener breakdown contribution at lower pulse energies, possibly alongside other processes such as above-threshold ionization. Fig. 1 b demonstrates excellent agreement () with data from the planar interface, showing PL intensity saturation around 8 nJ. This aligns with the Keldysh prediction of MPI dominance below 16 nJ. The photon energy for defect generation is times the photon energy from 790 nm laser, yielding eV for the planar interface and eV for the SIL interface, both of them are significantly lower than the displacement energies for Si (66 eV) and C (24 eV) in 4H-SiC, indicating defect formation mediated via laser-generated hot charge carriers rather than direct atomic displacement[75].
III Results
III.1 Characterisation of laser-written quantum emitters
The laser-written defects were characterized at room temperature with a home-built confocal microscope, as described in our earlier work [26]. A 780 nm diode laser provided continuous-wave optical excitation, and the emitted light passed through several long-pass filters at 800 nm before PL detection. By spatially mapping this PL emission over SILs that were laser-written into, we could identify laser-induced defects localized near the center of SILs. (Fig. 2).
For each SIL where a laser-written spot was detected, we performed measurements to determine the excitation power and brightness at which the emission saturates (power saturation measurement) and the normalized depth of the second-order correlation function at zero delay (), using low-jitter (average 40ps for the two channels) superconducting nanowire single-photon detectors (SNSPD, Single Quantum EOS). There is noticeable background luminescence from the surface (and edge of the SIL) but not within the SIL itself. This agrees with previous similar studies in 4H-SiC that notice intrinsic surface defects related to the surface oxide [55].
Stochastic distributions of emitting defects consistent with Poissonian statistics could be detected across the written region, with the number and brightness of defects increasing with increasing laser-writing power. To minimize the creation of multiple emitters, writing must ideally be performed in the regime where the expectation value of creating any emitters is much less than one; for example, for a Poissonian expectation value of , the probability of creating a single emitter is , while the probability of creating two or more emitters is . So, even for such a low yield, two emitters would be created approximately 5% of the time. In the case of the region with laser writing pulse energy PE=1.8 nJ shown in Figure 2, 9 written defects occur within a set of 30 SILs, including five confirmed single photon emitters and two defects consistent with two emitters, leading to a Poissonian expectation value of approximately .
The resulting values are reported in Table 1; we show the second-order correlation results before and after background subtraction. Fabrication using laser pulse energy 1.8 nJ created the most amount of emitters with , which is also illustrated in Figure 2. The SIL ID in Table 1 is with respect to the column and row of the SIL on the sample. For example, T30 is on column T and row 30 of the SILs processed, the characterisation for the spot in this SIL is also shown in Figure 3. The photoluminescence counts at saturation are also shown for each of these emitters.
We classify each PL spot based on the value to be either a single emitter (), multiple emitters , or as a PL spot not in the single-emitter regime . Therefore, as we decrease the laser writing power, we expect the likelihood of generating an emitter with () to increase so that there is an optimal laser writing power to generate single emitters reliably. This shows how narrow the window for effective fabrication of single color centers is using laser writing in SILs structures.
Throughout several months of measurements, we have seen no evidence of photo-bleaching, and the PL emission for all examined color centers have remained stable and reproducible.
| SIL ID | PE (nJ) | PL (kcps) | g (raw) | g (bkg. subtr.) |
| T30 | 1.9 | |||
| X18 | 1.8 | |||
| X23 | 1.8 | |||
| X26 | 1.8 | |||
| AA32 | 1.6 | |||
| Y17 | 1.8 | |||
| X20 | 1.8 |
To confirm the nature of the created quantum emitters, we examined their PL spectrum at low-temperature to identify characteristic zero-phonon lines (ZPL). The sample was placed in a cryostat at T= 4 K (Montana s100 Cryostation) and probed with a custom confocal setup (see Cilibrizzi et al. [8]). The emitters are excited with a 780 nm CW laser, and the PL spectra are measured with a grating spectrometer (OceanOptics QE Pro)). All retrieved spectra are shown in Supplementary Figure S5. Remarkably, we observe a wide spread of ZPL wavelengths in the spectral region 858nm-985nm. Out of 39 emission defects examined, 6 can have spectral lines that can be classified as silicon-vacancy (VSi) centers; these were, in particular, 1 V1’ center (ZPL at 858 nm), 4 V1 centers (h-site, ZPL at 861 nm) and 1 V2 center (k-site, ZPL at 916 nm). The remaining defects exhibit ZPLs at different wavelengths across this 858-985 nm range. These lines do not appear to match other unidentified lines previously reported in this wavelength range [76, 77, 78], but do fall within the theoretically-predicted range of VSi centers modified by nearby carbon anti-sites [79] and experimentally-observed range for silicon vacancies in etched membranes [80].
We further performed optically-detected magnetic resonance (ODMR) measurements at room temperature, in the frequency ranges associated with V1 (zero-field splitting 4 MHz) and V2 (zero-field splitting 70 MHz) centers. For as-written emission defects, no ODMR signals were observed in these ranges. However, after annealing the sample at 600∘C for 30 minutes in vacuum (mbar), a process known to improve the yield of VSi centers, we observed an ODMR signal associated with V2 for a single SIL (of 12 measured), though with a large linewidth of about MHz (see Fig S6 in Supplementary Information).
One possible explanation for the broad linewidths in the PL spectra and the lack of ODMR signal might be that the femtosecond laser creates other defects that are not optically active, in addition to the VSi. It is for example known, from molecular dynamics simulations [75] and photoluminescence and deep-level transient spectroscopy (DLTS) experiments [81, 82, 83], that carbon atoms feature much lower displacement energy than silicon atoms in SiC, so that the creation of e.g. carbon vacancies (V) is favored. The presence of charge traps, related to carbon vacancies and other defects [84, 60, 61], is expected to create electric and magnetic noise that may broaden the optical and magnetic resonance linewidths, respectively. Further experiments, beyond the scope of this manuscript, are needed to clarify the physics of femtosecond laser generation of point defects in SiC and optimize the laser writing parameters to improve the quality of generated quantum emitters.
III.2 Properties of single laser-written within a SIL
In Figure 3, the key properties are given of a single color center generated in SIL T30 with a laser writing pulse energy . The photoluminescence map (Fig. 3 a) shows a spot centered well with respect to the SIL profile (dark circle). The photoluminescent background outside of the SIL region corresponds to surface defects on the etched surface of the SiC, postulated to arise at the SiC-surface oxide interface [55, 28, 85, 86, 87]. Spectroscopy of the emitted light at cryogenic temperatures (T = 4K) yielded a ZPL at 861 nm, consistent with a V1-type VSi center. Furthermore, even without background subtraction, the normalized second-order correlation at zero delay , confirms that the emission comes from one single photon emitter. Taking power saturation measurements of the single VSi center in SIL T30 and a single electron-irradiated VSi center beneath a planar interface (Fig. 3, blue and green, respectively), the performance of the SIL-registered emitter could be benchmarked. Following our previous work [26], we determine an optical collection efficiency enhancement factor of 4.5 ( vs. ) and a power intensification factor of 9.1 ( vs. ), consistent with the upper levels of performance observed for single VSi centers generated randomly throughout the SIL by electron irradiation.
III.3 Registration of emitters to SILs
As mentioned earlier in Section II.1, laser writing through the SIL interface was carried out by manually aligning the laser focus to the center of the SIL at a depth of m, equivalent to the height of the lens structure. Contributions to the final emitter misalignment with respect to this target could be expected both from the random probability of generating a defect within the volume of highest writing intensity and the error in manual alignment, however these could be reduced by the focusing effect of the lens.
To quantify these distributions, spatial PL maps of SILs with emitter spots were characterized, using image analysis software (ImageJ) to extract the center of each SIL and emitter spot (further discussed in the Supplementary Information ). The absolute displacements of all measured emitters from their mean position are given in Fig. 4 a. The statistics follow a Rayleigh distribution with scale parameter m (dashed line, for reference).
When the absolute positions of the emitters with respect to the SIL are plotted (Fig. 4 b), clustering near the SIL center is evident, especially compared to emitter generation through electron irradiation in similar structures [26]. Focusing on the center region of the SIL (Fig. 4 c), it is apparent that the standard deviation of positions in the x-direction (nm) is higher than in the y-direction (nm). This is possibly due to the raster protocol for laser writing, where the writing proceeds along all columns in a row first before moving to the next row, so more steps are taken in x than in y.
Aligning to the SIL center can prevent the average position of generated defects from being at the center of the SIL due to systematic offsets from positioning or the writing angle. This is in contrast to writing in planar interfaces, where the locations of generated emitters are compared to the array of writing locations so that no mean offset is expected. This is also observed in Figure 4 c, where the mean emitter position is offset by 260 nm in x and 60 nm in y from the SIL center.
IV Conclusions
In this paper, we have demonstrated marker-free registration of single quantum emitters at the center of SILs, by femtosecond laser writing. The microlens both lower the laser dose required to generate the emitter and aid in positioning it near the center of the structure, with an accuracy of within in the y-direction and in the x-direction.
There remains an open question regarding the nature and diversity of created emitters and their performance for quantum technology applications. The formation of carbon-related defects, such as carbon vacancies, in SiC is known to be energetically quite favourable, suggesting that the laser irradiation might be creating other carbon-related defects in the immediate vicinity of the optically active VSi. ODMR properties have been measured, and none provided identifiable peaks. A further study after an anneal of the sample demonstrated that there are options to improve the spin properties of the defects. However, the yield of these results was low. Future studies will identify these defects and characterize their properties as a function of the laser treatment parameters.
In conclusion, this paper demonstrates the generation of single quantum emitters registered directly into microlens photonic structures in SiC, which can open new avenues for the scalable integration of spin-photon interfaces into quantum devices.
acknowledgments
We express our gratitude to Nguyen Tien Son and Ivan G. Ivanov for valuable discussions. We also thank Pasquale Cilibrizzi for his helpful guidance during the low-temperature measurements. This work is funded by the Royal Academy of Engineering (RF2122-21-129), the Engineering and Physical Sciences Research Council (EP/S000550/1, EP/V053779/1, EP/Z533208/1, EP/Z533191/1, EP/W025256/1), the European Commission (QuanTELCO, grant agreement No 862721), the Leverhulme Trust (RPG-2019-388) and Euramet (23NRM04 NoQTeS).
Supplementary Information
S0.1 Laser writing system
The laser writing system employed in this work is consistent with that described in a previous study [51]. The system comprises a Spectra Physics Mai Tai laser and a Spectra Physics Solstice amplifier, operating at a wavelength of 790 nm with a maximum repetition rate of 1 kHz. Pulse energy is regulated via a waveplate in combination with a Glan-laser polarizer. The beam is subsequently expanded before reaching the spatial light modulator (SLM, Hamamatsu Photonics X10468-02). A dichroic mirror directs the fabrication laser toward the sample while transmitting both the excitation and PL signals from an integrated room-temperature confocal system, featuring a 532 nm CW laser as the excitation source. Schematics of the laser writing system and confocal microscope are provided in Fig. S2 and Fig. S2, respectively.
S0.2 Description of the circle-fitting procedure
Here we describe the procedure to assess the position of the created emitter with respect to the SIL, based on PL maps with a stepsize of 0.13 m. First of all, we determine the SIL circular edge to retrieve its center. We used a data analysis application (ImageJ) to determine the SIL circular edges, using three different methods to verify the reproducibility of the outcomes. First, we use the fit circle tool to fit an exact circle around the SIL; second, we use the ellipse tool to fit precisely the SIL edge. Lastly, we used the profile tool to measure the changes in PL, which are expected to be high at the SIL edge as the surface of SIL itself fluoresces. We then fit a Gaussian to each of these PL peaks to pinpoint a precise position for the edges. Figure S3 shows a comparison between the three methods for the same SIL. The ellipse method in orange, the circle method in yellow and the blue lines indicate where the profiles would be taken to find the SIL edge as the PL counts rise relative to the background within the SIL.
Similarly, we compare results from two methods to determine the emitter position. First, we use the fit circle tool in ImageJ, we fit a circle around the emitter confocal spot, and retrieve its center. Second, we fit a Gaussian to the emitter confocal spot. Figure S3 shows the typical positions for these methods, where the blue lines again show the profile positions and the green shows the circle that would fit around the defect to determine the central point to be taken as the defect position.
We used the results for all SILs that generated a spot at 1.6 nJ writing power to benchmark the methods. The average SIL center using the ellipse and fitted circle methods were the same as , . The average calculated position relative to the SIL center using the profile method was in the x, and in the y. The average calculated position for the defect relative to the SIL center using the fitted circle’s method was in x and in the y-direction. Due to the defects being within the lens, we performed a correction to account for the magnification effects on the measured distance between the center of the spot and the SIL.
S0.3 XZ profiles in SILs
Fig. S4 presents a series of XY and XZ 2D PL images for defects fabricated with high pulse energy of 3.5 nJ (a) and low pulse energy of 2 nJ (b). These PL images indicate that the defect centers are registered near the center of the SIL in both lateral and axial directions, aligning well with the SIL’s focus. The elongated point spread function (PSF) observed in the axial PL profiles is attributed to the increased Abbe resolution in confocal microscopy. Quantitatively, the lateral and axial resolutions extracted from low pulse energy fabricated defect (110 nm and 170 nm, respectively) align with diffraction-limited expectations (100 nm lateral, 160 nm axial). In contrast, defect fabricated with higher pulse energy exhibits broadened emisssion profile (210 nm lateral and 985 nm axial, respectively), attributed to increased lattice damage induced by higher pulse energy that disrupts the point-source approximation.
S0.4 Optical Spectra of Emitters
To verify the characteristics of the quantum emitters produced, their photoluminescence (PL) spectrum was analyzed at low temperatures using a cryostat set to 4K (Montana s100 Cryostation) in combination with a custom confocal setup, as detailed by Cilibrizzi et al. [8]. The emitters were illuminated with a continuous wave (CW) laser at 780nm, and their PL spectra were recorded with a grating spectrometer (OceanOptics QE Pro) with a 800nm longpass filter. The spectra obtained are displayed in Supplementary Figure S3.
The analysis uncovered a broad distribution of zero-phonon lines (ZPL) within the 858nm-985nm spectral region. Out of 39 examined defect centers, six exhibited spectral lines typical of silicon-vacancy (VSi) centers: one V1’ center (ZPL at 858nm), four V1 centres (h-site, ZPL at 861nm), and one V2 center (k-site, ZPL at 916nm). Other defect centers showed ZPLs at various wavelengths within the 858-985nm range, which do not correspond to previously reported unidentified lines in this range but align with the theoretically predicted range of VSi centers modified by nearby carbon anti-sites.t
The broad line widths observed in the PL spectra and the lack of an ODMR signal might be attributed to the femtosecond laser’s creation of additional non-optically active defects alongside the V.
S0.5 ODMR after Anneal
An anneal at 600 degrees was conducted for 30 minutes on the sample in vacuum at mbar to investigate the stability of the observed defects, and whether they could be converted to silicon vacancies. We re-characterised 5 SILs, which originally showed no ODMR. One evidenced an ODMR peak after annealing, with 0.48% contrast and 22.36 MHz linewidth (figure S6).
The ODMR measurement is performed by focusing an off-resonant cw-laser (730 nm) on the a-plane side of the 4H-SiC sample inside the cryostat (Attocube Attodry 800) at 4K with a high-NA objective (Zeiss Epiplan-Neofluar 100x, NA 0.9). The spin state manipulation of the color center is orchestrated with the help of an Arbitrary Waveform Generator (QM-OPX) and a copper wire (50 diameter) running on top of the sample. The AWG feeds a radio-frequency signal with a power of -27dBm, amplified by a 44 dB amplifier (LZY-22+ Mini-circuits) and fed to the microwave antenna wire running over the sample. The manipulation of the spin states is realized once the frequency of microwave matches with the zero field splitting of ground state spin ( 70 MHz for a color center at = 0G). The difference in spin state population while sweeping the MW frequency is seen through the photoluminescence emitted by the V2 VSi center which is filtered using a dichroic mirror (Semrock FF925-Di01) and along pass filter (FELH950 Thorlabs). The emission is readout through a super conducting nanowire single photon detector (Single Quantum).